Find the values of all the trigonometric functions from the given information.
cos t =12/13, t is in quadrant I

sin t=
tan t=
cot t=
sec t=
csc t=

Answer by stanbon(75887)

Find the values of all the trigonometric functions from the given information.
cos t =12/13, t is in quadrant I
Since cos = x/r, x = 12 and r = 13
Then y = sqrt[13^2-12^2} = sqrt(25) = 5
==============
sin t= y/r = 5/13
tan t= y/x = 5/12
cot t= x/y = 12/5
sec t= r/x = 13/12
csc t= r/y = 13/5
=========================
Cheers,
Stan H.
-------------------------